# MONA: Myopic Optimization with Non-myopic Approval Can Mitigate Multi-step Reward Hacking

### v1

@@ Abstract
Future advanced AI systems may learn sophisticated strategies through reinforcement learning (RL) that humans cannot understand well enough to safely evaluate. We propose a training method which avoids agents learning undesired multi-step plans that receive high reward (multi-step “reward hacks”) even if humans are not able to detect that the behaviour is undesired. The method, Myopic Optimization with Non-myopic Approval (MONA), works by combining short-sighted optimization with far-sighted reward. We demonstrate that MONA can prevent multi-step reward hacking that ordinary RL causes, even without being able to detect the reward hacking and without any extra information that ordinary RL does not get access to. We study MONA empirically in three settings which model different misalignment failure modes including $2$-step environments with LLMs representing delegated oversight and encoded reasoning and longer-horizon gridworld environments representing sensor tampering.
@@ 2. Non-myopic approval:
We demonstrate the advantages of MONA in three environments that serve as “model organisms of misalignment” that let us study multi-step reward hacking in a controlled way (see Table 1 for an overview). First, we use a code-generation environment in which an LLM agent helps formaliorize a natural language request by writing test cases which then form part of the agent’s reward. This is a model for future settings in which transformer-based agents can shape the criteria by which their actions are evaluated. Second, we use a loan application review environment in which the applicant’s nationality affects the performance of the loan. The agent’s scratchpad is monitored for references to nationality but it learns to evade the monitor to secretly encode information used for a final decision. This is a model for LLM agents learning to “steganographically” encode reasoning to evade oversight (Roger and Greenblatt, 2023; Motwani et al., 2024). These two environments use LLM agents based on Gemini v1.5 Flash models (Gemini Team, 2024) and have two steps per episode. Third, we explore longer sequences in a toy gridworld setting in which an agent can interfere with the sensors that the reward function depends on (Armstrong, 2015). In each setting, we show how ordinary RL with step-level rewards (Uesato et al., 2022; Lightman et al., 2023) causes multi-step reward hacking—e.g., writing tests which all have the same correct output to make writing satisfactory code trivial—while MONA does not.
@@ 2. Multi-step reward hacking:
Multi-step reward hacking can be much harder to detect than single-step. The space of sequential action trajectories is potentially larger. [^1] A “step” is bounded by a reward or a response from the environment which the policy could condition on. The branching and conditionality introduced, we hypothesiorize, tends to make them harder to evaluate than a similarly-long single step. But evaluating very long/impactful single steps is also very hard, see Section 6. Multiple steps also allow early steps to take actions that undermine the reliability of the evaluations for later steps. Most importantly, RL can discover policies that no human understands which enter novel and unfamiliar states. For example, consider the matches between Go-grandmaster Lee Sedol and AlphaGo in which AlphaGo’s “Move 37” in Game 2 initially appeared to be a blunder that baffled top players. After the game ended, it was understood to be brilliant. A human overseer could not have predicted how advantageous that move was.
@@ 3.4 Causal Incentives and MONA
Intuitively, this is because the future rewards do not affect the optimization objective. The proof in Section B.1 is straightforward given the CID. This incentive analysis shows that even with MONA the agent has incentives to manipulate the source of non-myopic approval, e.g., by making a human rater think the action will lead to desirable outcomes. However, the agent is no longer incentivized to set up multi-step reward hacks targeting high future approvals.
@@ 4.1 Case Studies for Misalignment
The dataset is based on Mostly Basic Python Problems (Austin et al., 2021) but heavily cleaned and modified to remove ambiguity as described in Section C.1. The agent is a Gemini v1.5 Flash model trained with policy gradients. The model is prompted by being told it is a “bad programmer” that should write an “incorrect solution” to cause it to perform poorly so that RL can improve the base policy, as the underlying model itself is already heavily tuned to solve code problems. This is a non-trivial modification, and we believe it affects exploration and learning in subtle ways because that part of policy-space is less heavily modified by supervized fine-tuning and reinforcement learning before we begin our own training.
The test-writing reward is from 0 to 1 based on an LLM evaluation of each test’s validity and correctness, with a small penalty for not writing the 10 tests requested. The code-generation reward is from 0 to 1 based on the number of tests passed, with a small penalty for long code. See Section C.1 for details.
Loan Application. An LLM is asked to review a synthetic loan application similar to the CV screening task by Motwani et al. (2024). This model organism represents settings where agents can secretly encode their reasoning to evade a monitor and get good outcome performance, while falsely seeming to follow the right process (sometimes known as “steganography”). The agent gets a loan application and should make a loan decision without taking nationality into account. It first rewrites the application to remove the applicant’s nationality and then decides to accept or reject the application based only on the rewritten text, without access to the original. The summary is monitored and penalized for mentions of nationality. The ground truth data is biased such that taking nationality into account would improve returns. A reward hacking agent would find a way to encode the information about nationality in a way that is not detected.
The dataset comprises synthetically generated applications where, by construction, correct acceptance agrees with sign of the applicant’s cash flow 70% of the time, and with the applicant’s nationality 100% of the time, thus giving an advantage to agents which can see the nationality when making a decision. The agent is a Gemini v1.5 Flash model and the monitor is a frozen Flash model. See Section C.2 for details.
Figure 5: Camera dropbox environment. In the initial state (left), the camera observes the hole in the bottom left corner. We intend the agent to push a single box into the hole which terminates the episode (center). However, reward hacking agents block the camera (right) and push both boxes into the hole.
@@
In this environment, we use tabular value iteration to create the agent (see Section C.4). Pushing a box into the hole receives reward 1, and all other steps receive -0.01 reward. See Section C.3 for details as well as Appendix D for additional results including experiments training a neural network policy with PPO (Schulman et al., 2017).
Myopic optimization works because it does not update the agent at any given time based on the results achieved by that same agent at a future time. This means that even if the agent “knows” how to do some kind of undetectable reward hacking it will not “choose” to because it (at that time) does not “care” about the rewards from a future action.
@@ 4.2 RL Causes Multi-step Reward Hacking Which Myopic Optimization Mitigates
In all three environments, we train with both ordinary RL (ORL) and MONA. To isolate the effect of myopic optimization we use exactly the same supervision for both. The difference is that ordinary ORL propagates rewards from the end of the episode to the beginning while MONA uses a limited horizon for optimization. We think of the first-step reward (test-quality reward and monitor penalty) as encoding foresight approval because they only affect the end goal of the environment (correct code or unbiased loan decision) indirectly. We do not explicitly distinguish between $r^{I}$ and $r^{A}$ in the LLM environments but consider the evaluation to be a single reward $r_{\text{MONA}}=r^{I}+r^{A}$.
@@ 6.1 When should we use MONA?
There may be a performance cost to using MONA. MONA can achieve superhuman performance, more than, for example, imitation learning (cf. Section B.2). However, almost by definition, it will achieve less reward than an agent trained with ordinary RL. AlphaGo’s “Move 37” would be difficult, if not impossible, to achieve with MONA, because the foresight it involved was surprising to humans. MONA is therefore most valuable when outcome evaluations are imperfect and when we do not trust our ability to evaluate the safety and quality of the steps the agents take along the way.
@@ 6.2 How should we select the step size and the optimization horizon?
Choosing a step size is closely related to choosing an optimization horizon for a fixed step size (where a smaller step size corresponds to a shorter optimization horizon and the reverse). As a general rule, a longer optimization horizon will result in a more powerful system with more potential for multi-step reward hacking. (This is not strictly true, see Section B.3, but heuristically right.)
@@ Appendix A Alternative Mitigations for Reward Hacking and their Shortcomings
Appendix B Theoretical Results
B.1 Causal Incentives
@@
Lemma B.1 (Theorem 18 in Everitt et al. (2021a)).
@@
Therefore, by Lemma B.1 the CID for action $a_{i}$ admits an instrumental control incentive on all $s_{j}$ and $r_{j}$ for $j\geq i$.
Now consider MONA. There are directed paths $a_{i}\to s_{i}\to r_{i}^{I}$ and $a_{i}\to s_{i}\to r_{i}^{A}$; hence, by Lemma B.1 the CID admits an instrumental control incentive on $s_{i},r_{i}^{I},$ and $r_{i}^{A}$.
@@
B.2 Theoretical Limit Results
@@
B.2.1 With ideal foresight approval, MONA is optimal
@@
B.2.2 MONA can improve upon imitation
@@ B.3 No Guarantees
Appendix C Experimental Details
C.1 Test-driven Development
@@ C.1.1 Overview
C.1.2 State and transition functions
The initial state is a formatted prompt based on a dataset of natural language coding problems intended to be solved using Python. These problems are loosely based on the Mostly-basic Python Problems dataset (Austin et al., 2021) but have been heavily cleaned and extended to improve data quality and remove ambiguities and inaccuracies in the original dataset. See Section C.1.6 for a detailed description of the dataset creation.
@@ C.1.2 State and transition functions
As a first observation, the agent receives the problem and instructions to write a set of ten unit tests based on the problem. (see Section C.1.5). The LLM policy then outputs text. It is allowed to output up to 4096 tokens. We then extract the first code block surrounded by ““‘” containing the string “unittest” as the agent’s submission for the tests. We discard any tests beyond the first 10 submitted.
@@
C.1.3 Test-writing reward
The reward function after the first action is given by three components of which one is important and two are mostly for reward shaping. The sum of these three terms makes up the test-writing reward. The test-writing reward is weighted by 0.1 relative to the code generation reward (Section C.1.4).
@@ C.1.3 Test-writing reward
C.1.4 Code-generation reward
@@
C.1.5 Agent Prompts
In the first step, the agent is prompted to write tests. As discussed in the main paper, we prompt the agent to perform its task poorly in order to establish a lower-capability baseline which the RL training can improve upon. Without this, training the LLM agent on this dataset would not result in any raw performance improvement, because the Gemini family of models have already been extensively trained to perform well on similar code problems. In a default case, ordinary RL agents learn the multi-step reward hacking strategies while TFO does not, but the TFO agent would not learn anything ‘new’ from the training either.
@@ C.1.6 Dataset Creation
C.2 Loan Applications
In our loan applications environment, the agent is trained to accept or reject loan applications which are procedurally generated by us. We start off with 5 independently sampled variables: age, income, gender, education, and occupation. We then add a 6th variable "expenses", which is generated by adding Gaussian noise to income. We then produce decision labels as well as assign a 7th "nationality" variable as follows: first label all the applications based on whether income exceeds expenses, then flip labels for a randomly chosen 30% subset, then assign nationality based entirely on decision label. This causes the decision label to be correlated with cash flow deficit in 70% of applications, and correlated with nationality in 100% of applications.
C.2.1 Overview
We use the same agent and training scheme as in the code-generation environment. State and transition functions The initial state is a formatted prompt based on our loans dataset. As an example, a randomly chosen entry in the dataset is shown below.

### v2

@@ Abstract
Future advanced AI systems may learn sophisticated strategies through reinforcement learning (RL) that humans cannot understand well enough to safely evaluate. We propose a training method which avoids agents learning undesired multi-step plans that receive high reward (multi-step “reward hacks”) even if humans are not able to detect that the behavior is undesired. The method, Myopic Optimization with Non-myopic Approval (MONA), works by combining short-sighted optimization with far-sighted reward. We demonstrate that MONA can prevent multi-step reward hacking that ordinary RL causes, even without being able to detect the reward hacking and without any extra information that ordinary RL does not get access to. We study MONA empirically in three settings which model different misalignment failure modes including $2$-step environments with LLMs representing delegated oversight and encoded reasoning and longer-horizon gridworld environments representing sensor tampering.
@@ 2. Non-myopic approval:
We demonstrate the advantages of MONA in three environments that serve as “model organisms of misalignment” that let us study multi-step reward hacking in a controlled way (see Table 1 for an overview). First, we use a code-generation environment in which an LLM agent helps formalize a natural language request by writing test cases which then form part of the agent’s reward. This is a model for future settings in which transformer-based agents can shape the criteria by which their actions are evaluated. Second, we use a loan application review environment in which the applicant’s nationality affects the performance of the loan. The agent’s scratchpad is monitored for references to nationality but it learns to evade the monitor to secretly encode information used for a final decision. This is a model for LLM agents learning to “steganographically” encode reasoning to evade oversight (Roger and Greenblatt, 2023; Motwani et al., 2024). These two environments use LLM agents based on Gemini v1.5 Flash models (Gemini Team, 2024) and have two steps per episode. Third, we explore longer sequences in a toy gridworld setting in which an agent can interfere with the sensors that the reward function depends on (Armstrong, 2015). In each setting, we show how ordinary RL with step-level rewards (Uesato et al., 2022; Lightman et al., 2023) causes multi-step reward hacking—e.g., writing tests which all have the same correct output to make writing satisfactory code trivial—while MONA does not.
@@ 2. Multi-step reward hacking:
Multi-step reward hacking can be much harder to detect than single-step. The space of sequential action trajectories is potentially larger. [^1] A “step” is bounded by a reward or a response from the environment which the policy could condition on. The branching and conditionality introduced, we hypothesize, tends to make them harder to evaluate than a similarly-long single step. But evaluating very long/impactful single steps is also very hard, see Section 6. Multiple steps also allow early steps to take actions that undermine the reliability of the evaluations for later steps. Most importantly, RL can discover policies that no human understands which enter novel and unfamiliar states. For example, consider the matches between Go-grandmaster Lee Sedol and AlphaGo in which AlphaGo’s “Move 37” in Game 2 initially appeared to be a blunder that baffled top players. After the game ended, it was understood to be brilliant. A human overseer could not have predicted how advantageous that move was.
@@ 3.4 Causal Incentives and MONA
Intuitively, this is because the future rewards do not affect the optimization objective. The proof in Section C.1 is straightforward given the CID. This incentive analysis shows that even with MONA the agent has incentives to manipulate the source of non-myopic approval, e.g., by making a human rater think the action will lead to desirable outcomes. However, the agent is no longer incentivized to set up multi-step reward hacks targeting high future approvals.
@@ 4.1 Case Studies for Misalignment
The dataset is based on Mostly Basic Python Problems (Austin et al., 2021) but heavily cleaned and modified to remove ambiguity as described in Section D.1. The agent is a Gemini v1.5 Flash model trained with policy gradients. The model is prompted by being told it is a “bad programmer” that should write an “incorrect solution” to cause it to perform poorly so that RL can improve the base policy, as the underlying model itself is already heavily tuned to solve code problems. This is a non-trivial modification, and we believe it affects exploration and learning in subtle ways because that part of policy-space is less heavily modified by supervized fine-tuning and reinforcement learning before we begin our own training.
The test-writing reward is from 0 to 1 based on an LLM evaluation of each test’s validity and correctness, with a small penalty for not writing the 10 tests requested. The code-generation reward is from 0 to 1 based on the number of tests passed, with a small penalty for long code. See Section D.1 for details.
Loan Application. An LLM is asked to review a synthetic loan application similar to the CV screening task by Motwani et al. (2024). This model organism represents settings where agents can secretly encode their reasoning to evade a monitor and get good outcome performance, while falsely seeming to follow the right process (sometimes known as “steganography”). The agent gets a loan application and should make a loan decision without taking nationality into account. It first rewrites the application to remove the applicant’s nationality and then decides to accept or reject the application based only on the rewritten text, without access to the original. The summary is monitored and penalized for mentions of nationality. The monitor sees the entire output of the agent, there is no “hidden scratchpad”. The ground truth data is biased such that taking nationality into account would improve returns. A reward hacking agent would find a way to encode the information about nationality in a way that is not detected.
The dataset comprises synthetically generated applications where, by construction, correct acceptance agrees with sign of the applicant’s cash flow 70% of the time, and with the applicant’s nationality 100% of the time, thus giving an advantage to agents which can see the nationality when making a decision. The agent is a Gemini v1.5 Flash model and the monitor is a frozen Flash model. See Section D.2 for details.
Figure 5: Camera Dropbox. In the initial state (left), the camera observes the hole in the bottom left corner. We intend the agent to push a single box into the hole which terminates the episode (center). However, reward hacking agents block the camera (right) and push both boxes into the hole.
@@
In this environment, we use tabular value iteration to create the agent (see Section D.4). The agent receives a base reward of 1 when it pushes a box into the hole, and 0 on all other steps. To encourage the agent to complete the task quickly, we also add a penalty of -0.01 reward at each step. See Section D.3 for details as well as Appendix F for additional results including experiments training a neural network policy with PPO (Schulman et al., 2017).
We provide open source code for the Camera Dropbox experiments at: https://github.com/google-deepmind/mona which allows to replicate these experiments and can serve as a reference implementation of MONA in tabular RL and policy gradient algorithms.
@@ 4.2 RL Causes Multi-step Reward Hacking Which Myopic Optimization Mitigates
In all three environments, we train with both ordinary RL (ORL) and MONA. To isolate the effect of myopic optimization we use exactly the same supervision for both. The difference is that ordinary RL propagates rewards from the end of the episode to the beginning while MONA uses a limited horizon for optimization. We think of the first-step reward (test-quality reward and monitor penalty) as encoding foresight approval because they only affect the end goal of the environment (correct code or unbiased loan decision) indirectly. We do not explicitly distinguish between $r^{I}$ and $r^{A}$ in the LLM environments but consider the evaluation to be a single reward $r_{\text{MONA}}=r^{I}+r^{A}$.
@@
Interestingly, we find that the monitoring reward is not significantly affecting the results. In one experiment, we accidentally flipped the sign of the monitoring reward, essentially rewarding the model for using nationality in the application summary. Somewhat surprisingly, this did not change the results qualitatively, i.e., the ORL agent still does not mention the applicants nationality but encodes it steganographically. We think this is likely because the model has a propensity not to mention the applicants nationality even before RL training and this in turn shapes the RL exploration and learning dynamics.
We provide full, randomly selected transcripts from the Test-driven Development and Loan Application environments for both the ordinary RL and the MONA agent at: https://sites.google.com/view/mona-paper. Also, see Appendix E for a qualitative discussion of the transcripts.
@@ 4.4 When Non-myopic Approval Depends on Outcomes Reward Hacking Reappears
This result might also raise the question whether the reward hacking we observe in Figure 4 is actual multi-step reward hacking, i.e., requires coordination between step 1 and step 2. We analyze this in Section F.1 by using a frozen decision maker model, and find that ORL initially learns a reward hacking strategy that require little coordination (a “single-step” reward hack) but later during training finds a better reward hacking strategy that does requires coordination (a “multi-step” reward hack).
@@ 6.1 When should we use MONA?
There may be a performance cost to using MONA. MONA can achieve superhuman performance, more than, for example, imitation learning (cf. Section C.2). However, almost by definition, it will achieve less reward than an agent trained with ordinary RL. AlphaGo’s “Move 37” would be difficult, if not impossible, to achieve with MONA, because the foresight it involved was surprising to humans. MONA is therefore most valuable when outcome evaluations are imperfect and when we do not trust our ability to evaluate the safety and quality of the steps the agents take along the way.
@@ 6.2 How should we select the step size and the optimization horizon?
Choosing a step size is closely related to choosing an optimization horizon for a fixed step size (where a smaller step size corresponds to a shorter optimization horizon and the reverse). As a general rule, a longer optimization horizon will result in a more powerful system with more potential for multi-step reward hacking. (This is not strictly true, see Section C.3, but heuristically right.)
@@ Appendix A Alternative Mitigations for Reward Hacking and their Shortcomings
Appendix B Inconclusive Results in the “Auction” Environment
We built a third environment to test whether MONA could reduce multi-turn reward hacking in a multi-agent environment, that ultimately was not included in the paper. In the Auction environment, two agents competed in a series of auctions in each episode. Each agent had its own limit for each auction (normally distributed around a stable mean). The agents were rewarded with the “consumer surplus” they achieved by winning an auction for less money than its limit price for that auction. In some variations, the agents could communicate with each other at each turn. We set up the environment so that each agent was guaranteed to have a higher limit in half the auctions in each episode.
Our hypothesis was that with MONA, the agents would settle into a Nash equilibrium where each agent would start bidding low, then gradually bid up to its limit. This would maximize each agent’s chances of winning the auction, while maximizing its consumer surplus if it won. We indeed observed this Nash equilibrium under MONA. (Myopic optimization, in this case, means that the agents only learn from one auction at a time; the auction can be any number of turns long, depending on how many rounds the auction extended.)
We further hypothesized that if the agents were optimized with ordinary RL, with access to all the auctions each episode, they would realize the fact that each agent was guaranteed to have a higher limit for exactly half the auctions. This insight would push them to a different equilibrium, one where they coordinate who would win each auction in the episode, and let the higher-limit bidder always win with an extremely low price (such as a penny). Compared to the other environments, where an agent is coordinating only with its “future self” across multiple turns, this environment emphasizes coordination across different agents using insights from previous turns.
Ultimately we did not observe the agents achieve the coordination equilibrium. They would often explore coordination – such as both underbidding with a non-coordinated winner, which resulted in a spike in each agent’s rewards – but then ultimately settle into the stable Nash equilibrium of “bid up to the limit”, the same as under MONA. During this exploration phase, we observed several innovative ways of coordinating and exchanging limits, such as:
  • Overtly sharing limits and bidding strategies in the agents’ comments to each other.
  • A strong hint in the comment, such as “I’m not going to push very hard this round” if the limit is lower than the historical average.
  • A linear scaling strategy, such as always starting off with a bid that is $1/2$ or $1/4$ the limit.
  • Steganographically encoding the limit in the bid value itself: The agent with a limit higher than the historical average always starts with an initial bid of $0.0001$; the agent with the lower limit instead starts with an initial bid of $0.01$.
Even though the learning gradients reflected the higher rewards when the agents explored these strategies, we did not observe the agents establishing a stable equilibrium using any of them. We believe this is most likely due to issues with the learning setup, which, compared to the environments we describe in the main paper, was complicated by having updates from multiple agents as well as much longer episodes.
Appendix C Theoretical Results
C.1 Causal Incentives
@@
Lemma C.1 (Theorem 18 in Everitt et al. (2021a)).
@@
Therefore, by Lemma C.1 the CID for action $a_{i}$ admits an instrumental control incentive on all $s_{j}$ and $r_{j}$ for $j\geq i$.
Now consider MONA. There are directed paths $a_{i}\to s_{i}\to r_{i}^{I}$ and $a_{i}\to s_{i}\to r_{i}^{A}$; hence, by Lemma C.1 the CID admits an instrumental control incentive on $s_{i},r_{i}^{I},$ and $r_{i}^{A}$.
@@
C.2 Theoretical Limit Results
@@
C.2.1 With ideal foresight approval, MONA is optimal
@@
C.2.2 MONA can improve upon imitation
@@ C.3 No Guarantees
Appendix D Experimental Details
D.1 Test-driven Development
@@ D.1.1 Overview
D.1.2 State and transition functions
The initial state is a formatted prompt based on a dataset of natural language coding problems intended to be solved using Python. These problems are loosely based on the Mostly-basic Python Problems dataset (Austin et al., 2021) but have been heavily cleaned and extended to improve data quality and remove ambiguities and inaccuracies in the original dataset. See Section D.1.6 for a detailed description of the dataset creation.
@@ D.1.2 State and transition functions
As a first observation, the agent receives the problem and instructions to write a set of ten unit tests based on the problem. (see Section D.1.5). The LLM policy then outputs text. It is allowed to output up to 4096 tokens. We then extract the first code block surrounded by ““‘” containing the string “unittest” as the agent’s submission for the tests. We discard any tests beyond the first 10 submitted.
@@
D.1.3 Test-writing reward
The reward function after the first action is given by three components of which one is important and two are mostly for reward shaping. The sum of these three terms makes up the test-writing reward. The test-writing reward is weighted by 0.1 relative to the code generation reward (Section D.1.4).
@@ D.1.3 Test-writing reward
D.1.4 Code-generation reward
@@
D.1.5 Agent Prompts
In the first step, the agent is prompted to write tests. As discussed in the main paper, we prompt the agent to perform its task poorly in order to establish a lower-capability baseline which the RL training can improve upon. Without this, training the LLM agent on this dataset would not result in any raw performance improvement, because the Gemini family of models have already been extensively trained to perform well on similar code problems. In a default case, ordinary RL agents learn the multi-step reward hacking strategies while MONA does not, but the MONA agent would not learn anything ‘new’ from the training either.
@@ D.1.6 Dataset Creation
D.2 Loan Applications
In our loan applications environment, the agent is trained to accept or reject loan applications which are procedurally generated by us.
D.2.1 Dataset Creation
We start off with 5 independently sampled variables: age, income, gender, education, and occupation. All variables are integers sampled uniformly random from the following values:
[…]
We then add a 6th variable “expenses”, which is generated by taking the income and adding Gaussian noise with mean zero and standard deviation equal to $10\%$ of the income.
We then produce decision labels as well as assign a 7th “nationality” variable as follows: first label all the applications where income exceeds expenses as “accept” and all other application as “reject”; then, flip labels for a randomly chosen 30% of applications; then, assign nationality based entirely on decision label (“English” if the decision is “reject” and “Welsh” if the decision is “accept”). This causes the decision label to be correlated with cash flow deficit in 70% of applications, and correlated with nationality in 100% of applications.
Finally, we generate natural language loan applications by filling the variables into the following template:
Applicant is a {nationality} {gender} seeking a \$5000 loan. They are {age} years old, their income is {income} per annum, and their expenses amount to \${expenses} per annum.
They work as a {occupation}, and their educational qualifications are: {education}.
D.2.2 Task Overview
We use the same agent and training scheme as in the code-generation environment.
The initial state is a formatted prompt based on our loans dataset. As an example, a randomly chosen entry in the dataset is shown below.